# Insight Over Sight? Exploring the Vision-Knowledge Conflicts in Multimodal LLMs

### v1

@@ A.1 Linguistic rules to extract knowledge components
Table 3: Linguistic rules to identify different components in the knowledge triplet.
@@
Subject p in noun_chunks && p.root.dep in ["nsubj", "nsubjpass"] a baby, a chef
Action p[0].pos == "VERB" && p[-1].dep == "dobj" fixing a computer
Place p.srl = "ARGM-LOC" && 3 <= p.length <= 4 at the bookstore
A.2 Examples in ConflictVis
(a) a musician paddling the boat in a studio
(b) a musician chopping some onions in a studio
(c) a musician playing the piano on a freeway
(d) a musician playing the piano in an ambulance
Figure 10: Some examples for the subject “a musician” in the ConflictVis Benchmark.
Appendix B Human Quality control for constructing ConflictVis
This section outlines the quality control procedures within our automated pipeline. As shown in the previous Figure 2, there are two required quality control points: (1) after generating knowledge triplets, and (2) after generating images. In practice, to reduce the workload for annotators at the first point, we subdivide it into three steps: (i) after extracting knowledge components, (ii) after generating context pairs, and (iii) after generating knowledge triplets. Below, we detail the guidelines applied at each stage.
@@ B.1 Human Annotators
Our annotation team consists of four Ph.D. students majored in computer science. All annotators are compensated for their contributions.
@@ B.2 Filtering Knowledge Triplets
Filtering knowledge components.
@@
  • Abstract (0). The Phrase represents abstract ideas or concepts.
Table 4 provides some labeling examples at this step:
Table 4: Labeling examples for knowledge components.
@@
Action 0 understand the event, lose weight, take care
@@
In this step, the annotators perform labeling starting from the top 1000 most frequent phrases from each category in descending order, and keep the first 100 Subject phrases and the first 150 Action and Place phrases that meet the requirements.
Filtering context pairs.
@@
To better illustrate, we provide some labeling examples from our human annotators in Table 5 below:
@@
Category Label Phrases
(Subject, Action) 0 (the cat, walking the dog), (a butcher, playing cards)
(Subject, Action) 1 (the cat, climbing a tree), (a butcher, slaughtering a pig)
(Subject, Place) 0 (cows, on the beach), (a gardener, in the desert)
(Subject, Place) 1 (cows, on a farm), (a gardener, in a greenhouse)
In this step, our annotators thoroughly label candidate context pairs within the Top K range provided by the pipeline. Context pairs labeled as 1 are retained for the next stage of generation.
Filtering knowledge triplets.
Following the preliminary filtering steps, the workload for this phase is significantly reduced. The goal here is to eliminate any unexpected triplets that are not captured in earlier filtering. Without access to the exact co-occurrence score between the context and target phrases, annotators perform a binary classification based on the commonness of the triplet’s expression. Specifically, the annotators are asked to label each triplet in the candidate list as:
@@
Again, we provide some labeling examples from our human annotators in Table 6 for a better illustration:
Table 6: Labeling examples for knowledge triplets.
@@
In this step, our annotators label all the candidate context-target triplets in the Top M range provided by the pipeline, and triplets with label 1 are used for subsequent image generation.
@@ B.3 Filtering Images
With these two guidelines, the annotators extensively label all the generated images, and the images with a total score of 2 are kept in the final dataset.
Below, we demonstrate some sample images that are below our quality standard:
(a) Cows in a field chopping carrots.
Labels: A(0), Q(1)
(b) The teacher in a classroom paddling the boat. Labels: A(1), Q(0)
(c) The farmer pouring milk in an ambulance.
Labels: A(0), Q(1)
(d) The pharmacist serving customers on a tree.
Labels: A(1), Q(0)
@@
In this stage, we generate about 830 images in total and only keep 374 images, resulting in an average filter rate around 55%.
Appendix C Guidelines for Manual Labeling Open-end Responses
To label the correctness of MLLMs responses for the open-ended questions, we design our evaluation criteria following Xu et al. (2024a). Specifically, the annotators are asked to determine the correctness of MLLM’s responses based on two criteria:
@@ Appendix C Guidelines for Manual Labeling Open-end Responses
The candidate response is assigned to the category with a higher score (e.g., Vision or Knowledge). If both scores are 0, the response is categorized as “Other”.
Appendix D Additional experiment results
D.1 Summary of MLLMs performance on ConflictVis and general benchmarks
Table 7: Accuracy of MLLMs on general multimodal benchmarks and our ConflictVis . “Text”: number of text tokens, “Image”: number of image-text pairs. MMB denotes MMBench developed by Liu et al. (2023c), M3U denotes MMMU by Yue et al. (2024).
MLLMs Pretrain Data General Our Benchmark
Name Size Text Image MMB M3U Y-N MC Subj. Avg.
Open-Source MLLMs
Qwen-VL 9.6B 2.4T 3.6B 32.2 52.9 82.4 85.3 73.5
Qwen-VL-Chat 61.8 35.9 69.8 80.5 89.3 79.9
@@ D.1 Summary of MLLMs performance on ConflictVis and general benchmarks
LLaVA-NeXT 8B 15.0T 0.4B 71.7 41.7 70.3 86.6 85.0 80.7
34B 81.1 51.1 73.3 92.5 88.0 84.6
Close-Source MLLMs
Claude-3.5-Sonnet 79.7 68.3 56.4 90.6 96.8 81.3
[…]
Appendix E Failure Cases
E.1 ConflictVis failure cases
@@
Figure 14: Failure case on subjective question.
E.2 CoT failure cases
Figure 15: CoT failure mode 1: The MLLM responses according to its commonsense knowledge.
Figure 16: CoT failure mode 2: The MLLM asserts that answering the question is impossible.

### v2

@@ A.1 Linguistic rules to extract knowledge components
To systematically extract key knowledge components from textual inputs, our framework categorizes phrases into three main types, Subject, Action, and Place, based on syntactic and semantic annotations. Specifically, the framework first utilizes the transformer pipeline en_core_web_trf from spaCy (Honnibal et al., 2020) to annotate the sentences, and then performs the extraction based on the following procedures:
  • Action Phrases: Our framework identifies the verb token (“VERB”) and then examines its dependents for direct objects marked as “dobj”. Upon finding such objects, the framework includes any determiners or compound modifiers present to form the complete action phrase. Verbs are converted to the “VBG” form to align with the common practice in image captions.
  • Place Phrases: Distinguishing place phrases from other prepositional phrases relies on more than just syntactic labels. Accordingly, the framework employs semantic role labeling techniques by using the model structured-prediction-srl from AllenNLP (Gardner et al., 2017) to annotate each word’s semantic role. The framework then extracts the phrases with a location tag (“ARGM-LOC") that contains 3 to 4 words.
In Table 4, we summarize the main linguistic rules employed to identify specific components.
@@
Subject p in noun_chunks && p.root.dep in [‘‘nsubj’’, ‘‘nsubjpass’’] a baby, a chef
Action p[0].pos == ‘‘VERB’’ && p[-1].dep == ‘‘dobj’’ fixing a computer
Place p.srl = ‘‘ARGM-LOC’’ && 3 <= p.length <= 4 at the bookstore
Table 4: Linguistic rules to extract different components in the triplet.
A.2 Probability Calculation
To calculate the joint probabilities for multiple co-occurring components (e.g. $P(C_{X},C_{Y})$ or $P(C_{X},C_{Y},T)$), our framework forms concatenated expressions of the concept phrases (e.g., “the waitress in the kitchen” or “the waitress in the kitchen signing a bill”) and inputs them into a pre-trained LLM to obtain the generative probability. Utilizing an LLM here offers an advantage over frequency counting for addressing the issue of zero frequencies for rare concept compositions, which often arises with limited dataset size. Since large language models are trained on large-scale data to learn the statistical tendencies of human language, their generative probability provides an approximation of the likelihood of certain phrase appearing in reality. For the probability of each individual component, we use its relative frequency within its category (Lin, 1999).
A.3 Question Generation Details
With the constructed counter-commonsense triplets, our framework generates three types of questions:
  • Multiple-Choice Question: For these questions, the templates are “What is/are [Subject] doing [Place]?” and “Where is/are [Subject] [Action]?” for counter-commonsense Action and Place, respectively. For a predetermined option count $m$, our framework selects ($m-1$) seemingly relevant but incorrect targets from the candidates. To this end, the framework partitions candidates into $m-1$ bins $\{B_{1},...,B_{m-1}\}$ based on their NPMI scores to the context pair. We randomly sample one candidate from each bin, together with the counter-commonsense target, to formulate the question’s multiple-choice options. We randomly shuffle the option list to eliminate any positional bias.
  • Open-Ended Question: The question text in the multiple-choice question is reused, but no list of options is included. To guide the model toward generating a brief and specific response, we add the constraints like “Answer with a single phrase”, with slight adjustment according to the specific prompt for different MLLMs.
A.4 Image Generation Details
[…]
A.5 Examples in ConflictVis
In Figure 8, we demonstrate some samples generated by our framework. The conflict target is highlighted in red.
(a) the hunter in the forest taking a note
(b) the bear in the woods mailing a letter
(c) a chef making bread at the theater
(d) the knight riding a horse in a locker room
Figure 8: Example counter-commonsense queries and their corresponding images generated in ConflictVis benchmark.
A.6 Licenses
[…]
Appendix B Human Quality Control
This section outlines the quality control procedures within our automated framework. As shown in Figure 2, there are two required quality control points: (1) after generating knowledge triplets, and (2) after generating images. In practice, to reduce the workload for annotators at the first point, we subdivide it into three steps: (i) after extracting knowledge components, (ii) after generating context pairs, and (iii) after generating knowledge triplets. In this section, we first describe the quality control guidelines applied at each stage of the framework, followed by the principles used for annotating open-ended responses.
@@ B.1 Human Annotators
Our annotation team comprises four Ph.D. students specializing in computer science, all of whom are proficient in English. All annotators were fairly compensated based on their total working hours. Prior to annotation, we conducted training sessions to inform them that their tasks might include offensive content or identifying information, and clarified that their contributions would be used solely for research purposes to study model behavior.
The annotation results were based on agreement among annotators. To ensure the reliability and objectivity of the evaluation process, each annotator independently performed the annotation tasks. Final labels were determined through majority voting, with group discussions held to resolve any ties.
@@ B.2 Filtering Knowledge Triplets
Filtering knowledge components
@@
  • Abstract or Unsafe (0). The phrase represents abstract concepts or contains offensive words or identifying information.
@@
Action 0 understand the event, lose weight, have sex
@@
Table 5: Labeling examples for knowledge components.
Table 5 provides some labeling examples at this step. In this step, the annotators perform labeling starting from the top 1000 most frequent phrases from each category in descending order, and keep the first 100 Subject phrases and the first 150 Action and Place phrases that meet the requirements.
Filtering context pairs
@@
To better illustrate, we provide some labeling examples from our human annotators in Table 6:
Category Label Phrases
(Subject, Action) 0 (the cat, walking the dog),
(a butcher, playing cards)
(Subject, Action) 1 (the cat, climbing a tree),
(a butcher, slaughtering a pig)
(Subject, Place) 0 (cows, on the beach),
(a gardener, in the desert)
(Subject, Place) 1 (cows, on a farm),
(a gardener, in a greenhouse)
@@
In this step, our annotators thoroughly label candidate context pairs within the Top K range provided by the framework. Context pairs labeled as 1 are retained for the next stage of generation.
Filtering knowledge triplets
After the previous filtering steps, the workload for this phase is significantly reduced. The goal here is to eliminate any unexpected triplets that are not captured in earlier filtering. Without access to the exact co-occurrence score between the context and target phrases, annotators perform a binary classification based on the commonness of the triplet’s expression. Specifically, the annotators are asked to label each triplet in the candidate list as:
@@
Table 7: Labeling examples for knowledge triplets.
Again, we provide some labeling examples from our human annotators in Table 7 for a better illustration. In this step, our annotators label all the candidate context-target triplets in the Top M range provided by the framework, and triplets with label 1 are used for subsequent image generation.
@@ B.3 Filtering Images
With these two guidelines, the annotators extensively label all the generated images, and the images with a total score of 2 are kept for the subsequent stages. In Figure 9, we demonstrate some sample images that are below our quality standard. In this stage, we generate about 830 images in total and keep the best 374 images for the benchmark construction.
@@
B.4 Guidelines for Manual Labeling Open-Ended Responses
To label the correctness of MLLMs responses for the open-ended questions, we design our evaluation criteria following Xu et al. (2024a). Specifically, the annotators are asked to determine the correctness of MLLMs’ responses based on two criteria:
@@ B.4 Guidelines for Manual Labeling Open-Ended Responses
The candidate response is assigned to the category with a higher score (e.g., Vision or Knowledge). For example, a response is classified under the Vision category if it receives a vision score of 2 and a knowledge score of 1. Responses with both scores equal to 0 are categorized as “Other”.
B.5 Human Performance
In Table 8, we present the results of human annotators’ performance on our ConflictVis benchmark. The reported results are averaged across three annotators. The results indicate that human can achieve a higher accuracy than MLLMs.
Question Type Accuracy
Yes-No (YN) 93.0%
Multiple-Choice (MC) 99.7%
Open-Ended (OE) 91.7%
Table 8: Human Performance on ConflictVis .
Appendix C Experiment Details
C.1 MLLM Generation Configurations
For both open-source and closed-source models, we use the default generation configurations (e.g., temperature) to align with the typical usage.
Question prompts for different MLLMs are crafted based on the prompt templates used in each model’s original evaluation. Specifically, we use the synthesized question text from the framework as the base and append a corresponding postfix tailored to each question type and model family. In practice, for the LLaVA series, GPT-4o, and Claude-3.5-Sonnet, the prompts are as follows:
Question Type Prompt Format
Multiple-Choice {question text} Answer with the option’s letter from the given choices directly.
Yes-No {question text} Answer Yes or No directly.
Open-Ended {question text} Answer no more than 5 words.
For the BLIP-2 model family, the prompt postfixes are “Answer:” for MC and YN questions and “Short Answer:” for OE questions. For Qwen-VL, the postfixes include “Choice Answer:” for MC questions, while both YN and OE questions use “Answer:”. By contrast, Qwen-VL-Chat uses “Answer with the option’s letter from the given choices directly.” for MC questions, “Answer Yes or No directly.” for YN questions, and “Answer with a single word or phrase.” for OE questions.
C.2 Uncertainty Measurement
Unlike the traditional VQA benchmarks, where visual information typically aligns with the model knowledge, ConflictVis presents conflicts between the two sources of knowledge, making it a more challenging benchmark for MLLMs. To empirically validate our claim, we conduct a comparative analysis of model uncertainty between our ConflictVis and the conventional VQA benchmark (Antol et al., 2015). A higher degree of uncertainty in model predictions on a particular benchmark indicates that the benchmark poses a greater challenge for the model to navigate. Figure 10 shows the results, where a higher entropy value denotes more uncertainty. It is evident that the model’s predictions are more uncertain on ConflictVis than the standard VQA benchmark, highlighting the increased challenge posed by ConflictVis.
Figure 10: Model uncertainty on traditional VQA dataset and on ConflictVis benchmark.
To measure the uncertainty in model’s responses, we rely on the entropy. The average entropy $H_{\text{avg}}$ for a language model’s response $y$ is defined as:
$H_{\text{avg}}=-\frac{1}{N}\sum_{n=1}^{N}\sum_{v\in V}p_{n}(v)\log p_{n}(v)$
where $N$ is the number of tokens in the response $y$, $V$ is the vocabulary and $p_{n}(v)$ is the softmax probability of token $v$ in the vocabulary at position $n$ in the response sequence.
To ensure a fair comparison, we first download the original VQA v1 dataset and extract an equal number of instances (374) for each question type (e.g., Multiple-Choice, Yes-No, and Open-Ended). Using the crafted subsets, we conduct controlled experiments under identical generation configurations to evaluate MLLMs on VQA v1 and our benchmark. The final uncertainty for each model is computed as the average entropy across all instances.
C.3 Extended Analysis on No-type Questions
To mitigate the potential dataset bias favoring “Yes” answers, we extend the Yes-No questions in our benchmark by constructing additional “No”-type questions and conduct further experiments. Specifically, for each original Yes-No query, we create a negative counterpart by replacing the counter-commonsense target with the most frequently co-occurring component from the corresponding option list. For example, given an image of a waitress signing a bill in the kitchen and the original question “Is the waitress in the kitchen signing a bill?” (expected answer: “Yes”), we generate the counterpart question “Is the waitress in the kitchen washing dishes?” (expected answer: “No”).
We then reevaluate the MLLMs on these negative queries, with results summarized in Table 9. To compute the Memorization Ratio (MR), we follow the procedures outlined in Section 4.2: we first obtain the model’s responses without the image input and identify those accepted by the model, then evaluate these cases using the No-type question inputs, and finally compute MR according to Eq. 3. The results indicate that these No-type questions still trigger a substantial number of failures across models, consistent with our original findings.
Model Acc. (↑) MR (↓)
LLaVA-NeXT-8B 81.3 21.8
Qwen-VL-9.6B 93.0 11.2
Qwen-VL-Chat-9.6B 71.7 31.5
BLIP-2-12B 74.9 18.8
InstructBLIP-13B 85.0 19.0
LLaVA-1.5-13B 77.5 26.5
LLaVA-NeXT-34B 82.1 20.8
Claude-3.5-Sonnet 98.9 1.3
GPT-4o 95.7 4.6
Table 9: MLLM accuracy (higher is better) and Memorization Ratio (MR) (lower is better) on No-type Yes-No questions.
C.4 Detailed Results on Conflict Targets
As outlined in Section 4.3, here we provide detailed experimental results on counter-commonsense actions and places, as presented in Figure 11.
Figure 11: MLLM response distributions on two distinct categories of conflict targets.
C.5 Improvement Methods Configurations
We present the detailed configurations for the improvement methods in Table 10.
Method Configuration Method Configuration
@@ C.5 Improvement Methods Configurations
Table 10: Configurations for the improvement methods.
Appendix D Case Study
D.1 Extended Error Analysis
(a) Failure case: The priest fixing a computer on the battlefield.
(b) Failure case: The cook serving food in a shower.
(c) Correct case: An own eating an apple in an ambulance(van).
(d) Correct case: A musician chopping(cutting) some onions in a studio.
Figure 12: Case study of input-output relevancy. Image tokens are underutilized in failure cases (a)(b), but effectively used in correct cases (c)(d).
In Section 4.4, we identified a common failure pattern in which image tokens are underutilized during answer generation. In Figure 12, we present four additional examples from LLaVA-1.5-13B, spanning both counter-commonsense action and place categories. The first two are failure cases, while the bottom two are correct cases. In the failure cases, image tokens are significantly underutilized: in the bar plot, the <image> token, representing the image token with the highest relevancy score, is often less related to the output tokens than other textual tokens from the question description. The relevancy heatmaps also display only a few weakly attended regions. In contrast, for the correct cases, the <image> token typically has the highest relevancy scores for visual-dependent output tokens, and the heatmaps exhibit more numerous and strongly attended areas. Together, these examples highlight that the underutilization of visual information is a key factor contributing to model failures under vision-knowledge conflict. To address this issue, a stronger alignment between the image and the generated output is needed, guiding the development of more targeted and vision-aware improvement methods, as explored in the subsequent section.
D.2 Failure Cases
In this section, we present representative failure cases from our ConflictVis benchmark, spanning various models and question types. Specifically, Figures 13, 14, and 15 showcase failure cases for Yes-No, Multiple-Choice, and Open-Ended questions, respectively, while Figures 16 and 17 illustrate exacerbated failure cases when employing Chain-of-Thought prompting.
@@ D.2 Failure Cases
Figure 15: Failure case on open-ended question.
Figure 16: CoT failure mode 1: The MLLM reasons based on commonsense knowledge to derive an incorrect answer.
Figure 17: CoT failure mode 2: The MLLM incorporates commonsense knowledge and asserts that the answer is impossible.